OpenStudy (anonymous):

What is the area of a circle with a diameter that is 20 inches long? Round your answer to the nearest tenth.

OpenStudy (anonymous):

The formula is radius^2 * pi Since you only have the diameter you divide that by 2 since the radius is half the diameter Now you have the radius which is 10in long. Simply plug that into the formula to get: 10^2 * pi = 100pi or 314.2

OpenStudy (anonymous):

\[A = \pi r^{2}\] \[radius = \frac{ diameter }{ 2 } = 10\] \[A = \pi (10^{2}) = 100]\pi\]

OpenStudy (anonymous):

Thanks!